Session: Jun 2006 (session 2)
Wavelength: 5 cm


To give feedback on this experiment, click

Experiment Feedback

Medicina (G.Maccaferri at Fri Jun 9 14:02:24 2006) Success :

Hartebeesthoek (Jonathan Quick at Fri Jun 9 14:10:42 2006) Success :
No known problems

Torun (Kaz Borkowski at Fri Jun 9 14:25:49 2006) Success :
No known problems.

Onsala (Michael Lindqvist at Sat Jun 10 07:35:46 2006) Minor failures :
Fringe test said: No Onsala fringes from VC#03, upper and lower sidebands (LCP). However, VC03 looked OK. Otherwise no known problems.

Westerbork (Tony Foley at Mon Jun 12 06:46:58 2006) Success :
Single dish was used. Please change the coordinates in the vexfile accordingly. no known problems

Effelsberg (A. Polatidis at Mon Jun 12 08:44:59 2006) Success :

Noto (P. Cassaro at Sat Jun 17 10:47:11 2006) Success :

Jodrell_Bank (Alastair Gunn at Tue Jun 20 09:10:53 2006) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Tue Jun 20 09:11:36 2006) Failure : RECE
The Darnhall telescope was used. No fringes found due to bandswitch being in wrong position. Subsequently resolved. No other known problems.

EVN experiment feedback facility: entire session

Entire Session

Session: Jun 2006 (session 2)
Wavelengths: 30cm 18/21 5 6 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Jun 5 12:32:29 2006) Minor failures :
This note concerns the UHF experiments only. Shortly before the first experiment we could manage to install the Lange coupler to produce two circular polarizations instead of the linear polarization announced in the block schedule. Also, we failed to install the phase cal.

Jodrell_Bank (Alastair Gunn at Tue Jun 20 09:14:12 2006) Success : PCAL
The MkII telescope was used (Jb2). During this session the Mk2 phase cal. signal was found to be too high and could not be easily reduced in strength. I therefore switched off the phas cal. signal for the entire session (for Mk2 only).